# Interpretable global minima of deep ReLU neural networks on sequentially separable data

### v2

@@ Abstract.
1. Introduction
In this work, continuing the program started in [CE23a, CE23b], we provide an interpretation for the action of the layers of a neural network, and explicitly construct global minima for a classification task with well distributed data. This characterization is done in terms of a map which encodes the action of a layer given by a weight matrix $W$, a bias vector $b$, and activation function $\sigma$,
$\tau_{W,b}(x)=(W){}^{+}\left(\sigma(Wx+b)-b\right),$ (1.1)
where $(W){}^{+}$ denotes the generalized inverse of $W$. A neural network with $L$ layers can be seen as an application of $L-1$ such endomorphisms $\tau$ on input space, which curate the data, followed by an affine function in the last layer mapping to the output space,
$x^{(L)}$ $=W^{(L)}(x)^{(\tau,L-1)}+b^{(L)},$ (1.2)
where $x^{(\tau,L-1)}$ represents the $L-1$ consecutive applications of $\tau$ maps to an input $x$, and $W^{(L)},b^{(L)}$ are appropriate cumulative parameters. While this works for any given activation function, we specialize to the case of a ReLU neural network, and in this context call $\tau$ the truncation map. An incomplete, but highly intuitive visualization of its action comes from looking at cones: For any cone in input space, there exist $W$ and $b$ such that $\tau_{W,b}$ acts as the identity on the forward cone, and projects the entirety of the backward cone to the base point.
Consider data in $\mathbb{R}^{M}$ to be classified into $Q$ classes. In [CE23a], the tools outlined above were used to construct explicit global minimizers for a network with constant width $M=Q$ and “clustered” data. In this paper, we extend this to the case where the width of the hidden layers is allowed to decrease monotonically from $M$ to $Q$, for $Q\leq M$. As a corollary, we show a reduction in the total number of parameters needed to interpolate the data. Thus, our first result (informally stated) is
Theorem (Corollary to Proposition 4.1).
@@ 1. Introduction
Figure 1. A schematic representation of the action of a layer given by a truncation map. On the left, the original data with three classes contained in balls around the means $\overline{x_{0,j}}$, with $\mathcal{X}_{0,1}$ sitting in the backward cone $\mathfrak{C}_{\theta}$, and the remaining classes sitting in the forward cone. On the right, the truncation map has pushed all of $\mathcal{X}_{0,1}$ to a single point, the base of the cone. In green, we see the 2-simplex generated by the means, and the barycenter $\overline{x}$.
@@
As the angular aperture of a cone gets larger, it becomes a hyperplane. These are a well known classification tool, and there are algorithms dedicated to finding hyperplanes maximizing the distance to the two sets being separated, such as support vector machines (SVMs). Recent works show that under some circumstances, ReLU networks trained with gradient descent to perform binary classification tend to find max-margin hyperplanes, see e.g. [PL20]. [^1] The tendency of gradient-based algorithms to find critical points which generalize well or are simple in some way, e.g. low-rank, is called implicit bias or implicit regularization.
The same algorithms can be adapted to the case of multiple classes, by reducing to binary classification via several one-versus-all or one-versus-one strategies. However, a neural network tasked with performing multiclass classification must yield a single function capable of classifying. When the data is clustered as above, figure 1 shows how the layers of the network will push each cluster into a single point. By considering hyperplanes as limits of cones, this can be generalized to data for which at each step, one class gets mapped into a single point. Thus, each layer acts to find a one-versus-all hyperplane. If this can be done so that at the end there are $Q$ linearly independent points, the last layer will complete the classification with zero loss. We define data which satisfies this condition as sequentially linearly separable, and this leads to our next result (informally stated)
Theorem (Theorem 5.2).
$\left\{(\theta_{\ell},\nu_{\ell},\mu_{\ell})\right\}_{\ell=1}^{Q}\subset(0,\pi)\times\mathbb{R}^{M}\times(0,1),$ (1.3)
corresponding to triples of angles, normal vectors and a line segment of base points (resp.) describing cones and hyperplanes.
In [CE23a], the inductive proof of the main result was adapted to give $12^{Q}-$ 1 suboptimal local minima, associated to the binary choice for each cluster to have the truncation map act trivially or collapse the data to a point. However, the proof of minimality relied on an exact expression for the norm of local minima in the case $M=Q$ found in [CE23b, Theorem 3.2]. The assumption $M=Q$ is dropped for our results here, so the local minima cannot be found in the same way. Moreover, the sequential separability of the data in Theorem 5.2 can very highly dependant on the ordering chosen (cf. figure 6 in section 5), so that the truncation maps cannot be chosen independently.
Finally, we note that while we do not consider layers with arbitrary width, increasing the width would only add unnecessary parameters, as we assume the data to be suitably linearly separable. It is known that data which is not linearly separable can be classified by an SVM with the use of kernel machines ([BGV92]) which map the data into a space of higher dimension in such a way that in this new space, a separating hyperplane can be found. Moreover, neural networks approximate kernel machines in the infinite width limit ([LSDP+18, JGH18]). With these considerations, it could be the case that for data that is not linearly separable, the first layers of a neural network could be seen as describing a map to a higher dimensional space where the conditions for sequential linear separability are now satisfied, and the results above can be applied.
1.1. Related work
We give here a short and incomplete survey of related results.
Since the ReLU function acts as the identity when activated, it is instructive to compare with the linear case. The loss landscape of linear neural networks is very well understood. Its study started with the classic work [BH89], which proved (under some assumptions on the architecture and data) that a shallow (one hidden layer) linear neural network with square loss has only global minima and saddle points, and that every critical point is a projection of the least-squares estimator onto a subspace spanned by the eigenvectors of a relevant matrix determined by the data and label matrices. [^2] That is, a critical point with $rank(W_{2}W_{1})=r$ satisfies $W_{2}W_{1}=U_{r}U_{r}^{T}YX^{+}$, where $U_{r}$ is formed by $r$ columns of $U$, the matrix of eigenvectors of $\Sigma=Y(X){}^{+}XY^{T}$. See next section for notation. [Kaw16] extended this to deep linear networks, proving under milder conditions (both on data and architecture) that there are no suboptimal local minima, and giving some properties of the Hessian of the loss at saddle points. [ZL18] provided complete characterizations of all critical points of a deep linear network with square loss, arbitrary architecture and very mild assumptions on the data. Recently, [AMG22] gave a complete second order analysis of the loss landscape of deep linear networks, with the same assumptions as [Kaw16].
The loss landscape of ReLU neural networks, on the other hand, is not as well understood. Following assumptions adopted in [CHM+15], in [Kaw16] the activation functions are treated as random variables, and what is computed is the loss with expectation; under these assumptions, the loss reduces to the loss of a deep linear network. [Tia17] studied ReLU networks with no hidden layers, Gaussian distributed input data and target function a neural network with fixed weights, and found that certain types of critical points are not isolated. [ZL18] studied shallow ReLU networks with very mild assumptions on the data, characterized all critical points of the square loss for parameters in certain regions of parameter space (depending on the activation pattern of the neurons, and not covering the whole space), and gave an example of a network for which there is a suboptimal local minimum. While not studying critical points of a loss function directly, [GLMW22, GLR23] studied functional properties of ReLU neural networks, such as “functional dimension” [^3] Defined as the rank of the Jacobian of the network with respect to its parameters., the level sets of which are contained in level sets of any loss function.
Finally, the results here show how a trained neural network could take within-class variability to zero. This is one of the features of neural collapse, a phenomenon which has been empirically shown in [PHD20, HPD22] to emerge when training a neural network towards zero loss.
1.2. Outline of paper
In section 2, we clearly define the types of neural networks and the cost function under consideration and fix notation. In section 3, we define the truncation map, remark on its action in terms of cones in input space and cumulative parameters, and prove key lemmas. In section 4, we give precise statements for Proposition 4.1 and a corollary. In section 5, we define sequentially linearly separable data, and give a precise statement and proof for Theorem 5.2. In section 6, we prove Proposition 4.1. Finally, we collect some basic results on geometry of cones in appendix A.
2. Setting and notation
We will be dealing with neural networks of the form
$x^{(0)}$ $=x_{0}\in\mathbb{R}^{d_{0}}\,\,\text{ initial input,}$,
$x^{(\ell)}$ $=\sigma(W_{\ell}x^{(\ell-1)}+b_{\ell})\in\mathbb{R}^{d_{\ell}},\text{ for }\ell=1,\cdots,L-1,$ (2.1)
$x^{(L)}$ $=W_{L}x^{(L-1)}+b_{L}\in\mathbb{R}^{d_{L}},$
$X^{(\ell)}=\sigma\left(W_{\ell}X^{(\ell-1)}+B_{\ell}\right),$ (2.2)
where
$B_{\ell}:=b_{\ell}u_{N}^{T}\in\mathbb{R}^{d_{\ell}\times N}$ (2.3)
for
$u_{N}:=(1,\cdots,1)^{T}\in\mathbb{R}^{N}.$ (2.4)
$\overline{x_{0,j}}:=\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}x_{0,j,i}$ (2.5)
and the differences
$\Delta x_{0,j,i}:=x_{0,j,i}-\overline{x_{0,j}}.$ (2.6)
$\overline{X_{0,j}}$ $:=X_{0,j}-\Delta X_{0,j}$ $\in\mathbb{R}^{M\times N_{j}},$ (2.7)
$Y$ $:=\left[y_{1}\cdots y_{j}\cdots y_{Q}\right]$ $\in\mathbb{R}^{Q\times Q},$
Then $\mathcal{X}_{0,i}=\operatorname{Col}(X_{0,i})$, the set of columns of $X_{0,i}$, and similarly $\mathcal{X}_{0}=\operatorname{Col}(X_{0})$.
$(\overline{X_{0}^{red}}){}^{+}=(\overline{X_{0}^{red}}^{T}\overline{X_{0}^{red}})^{-1}\overline{X_{0}^{red}}^{T},$ (2.8)
$A(A){}^{+}A$ $=A,$
$(A){}^{+}A(A){}^{+}$ $=(A){}^{+},$ (2.9)
$A(A){}^{+}\text{ and }(A){}^{+}A$ are symmetric.
Recall that $(A){}^{+}A$ is an orthogonal projector to $(\ker(A))^{\perp}$, $A(A){}^{+}$ is an orthogonal projector to $\operatorname{ran\,}(A)$, and $(A){}^{+}$ is a left (resp. right) inverse of $A$ is injective (resp. surjective).
Finally, for an appropriate neural network defined as in (2) with $d_{0}=M$ and $d_{L}=Q$, we will be considering the square loss, given by the $\mathcal{L}^{2}$ Schatten (or Hilbert-Schmidt) norm,
$\left\lVert{A}\right\rVert_{\mathcal{L}^{2}}:=\sqrt{\operatorname{Tr}(AA^{T})},$ (2.10)
so that the cost function is
$=\frac{1}{2}\sum_{j=1}^{Q}\sum_{i=1}^{N_{j}}\left\lvert{x_{0,j,i}^{(L)}-y_{j}}\right\rvert^{2}.$ (2.11)
[…]
$\mathcal{C}[(W_{i},b_{i})_{i=1}^{L}]:=\mathcal{C}[X_{0},Y^{ext},(W_{i},b_{i})_{i=1}^{L}].$ (2.12)
Remark 2.1.
Note that another possible choice for the cost function would be to average over each class separately, so that
$=\left\lVert{X^{(L)}-Y^{ext}}\right\rVert_{\mathcal{L}_{\mathcal{N}}^{2}},$ (2.13)
where $\mathcal{N}:=\operatorname{diag}(N_{j}{\textbf{1}}_{N_{j}\times N_{j}}|j=1,\cdots,Q)$ and $\langle A,B\rangle_{\mathcal{L}_{\mathcal{N}}^{2}}:=\operatorname{Tr}(A\mathcal{N}^{-1}B^{T})$.
The cost $\mathcal{C_{\mathcal{N}}}$ can be rewritten as follows
$=\sum_{j=1}^{Q}\left(\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\left\lvert{x_{0,j,i}^{(L)}-\overline{x_{0,j}^{(L)}}}\right\rvert^{2}+\left\lvert{\overline{x_{0,j}^{(L)}}-y_{j}}\right\rvert^{2}\right)$ (2.14)
$=\sum_{j=1}^{Q}\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\left\lvert{W^{(L)}\Delta x_{0,j,i}^{(\tau,L-1)}}\right\rvert^{2}+\sum_{j=1}^{Q}\left\lvert{W^{(L)}\overline{x_{0,j}^{(\tau,L-1)}}+b^{(L)}-y_{j}}\right\rvert^{2},$
and since both terms are non-negative, the cost will be zero if, and only if, each term is zero. The global minimum with zero loss does not depend on the choice of weights for the cost function, [^4] That is, the constants in $\mathcal{N}$ could be any sequence $(\alpha_{j})_{j=1}^{Q}$; common choices are $\alpha_{j}=1$ or $\alpha_{j}=\frac{1}{2}$ for all $j=1,\cdots,Q$. and so we have the following interpretation for the results presented in this paper: The truncation maps implemented by the hidden layers have the role of minimizing the first term, by which all variances are reduced to zero, while the last (affine) layer minimizes the second term, whereby the class averages are matched with the reference outputs $y_{j}$.
3. Truncation map, cones and cumulative parameters
We recall here the definition of the truncation map from [CE23a], making the necessary modifications for the case where the width is not constant and the weight matrices might not be square.
Definition 3.1.
$\tau_{W,b}:\mathbb{R}^{M_{0}}$ $\to\mathbb{R}^{M_{0}}$
$x$ $\mapsto(W){}^{+}\left(\sigma(Wx+b)-b\right),$ (3.1)
where $(W){}^{+}$ is the generalized inverse defined as in (2).
$\tau_{W,b}(X)=(W){}^{+}\left(\sigma(WX+B)-B\right).$ (3.2)
where $B=b\,u_{N}^{T}$. This map encodes the action of a layer of a neural network and represents it in the domain. It has a nice recursive property which makes it possible to extend this representation to several layers. The following is a small generalization of the result from [CE23a] to the case where the dimension of the intermediate spaces given by the hidden layers is allowed to decrease.
Proposition 3.2.
Assume $M=d_{0}\geq d_{1}\geq\cdots\geq d_{L}=Q$, $X^{(\ell)}$ corresponds to a hidden layer of a neural network defined as in (2), and all of the associated weight matrices $W_{\ell}$ are full rank. Then the truncation map defined satisfies
$X^{(\ell)}=W_{\ell}\,\tau_{W_{\ell},b_{\ell}}(X^{(\ell-1)})+B_{\ell}.$ (3.3)
Moreover, defining the cumulative parameters
$W^{(\ell)}:=W_{\ell}\cdots W_{1},\quad\text{ for }\ell=1,\cdots,L,$ (3.4)
and
$b^{(\ell)}:=\begin{cases}W_{\ell}\cdots W_{2}b_{1}+W_{\ell}\cdots W_{3}b_{2}+\cdots+W_{\ell}b_{\ell-1}+b_{\ell},&\text{ if }\ell\geq 2,\\ b_{1},&\text{ if }\ell=1,\end{cases}$ (3.5)
we have
$X^{(\ell)}$ $=W^{(\ell)}(X_{0})^{(\tau,\ell)}+B^{(\ell)},$ (3.6)
for $\ell=1,\cdots,L-1$, and
$X^{(L)}$ $=W^{(L)}(X_{0})^{(\tau,L-1)}+B^{(L)},$ (3.7)
where
$=\tau_{W^{(\ell)},b^{(\ell)}}(X_{0})^{(\tau,\ell-1)}.$ (3.8)
Proof.
Using the fact that $W_{\ell}(W_{\ell}){}^{+}={\textbf{1}}_{d_{\ell}\times d_{\ell}}$ for all $\ell=1,\cdots,L$, we have
$X^{(\ell)}$ $=\sigma(W_{\ell}X^{(\ell-1)}+B_{\ell})$
$=W_{\ell}(W_{\ell}){}^{+}\sigma(W_{\ell}X^{(\ell-1)}+B_{\ell})$
$=W_{\ell}(W_{\ell}){}^{+}\left(\sigma(W_{\ell}X^{(\ell-1)}+B_{\ell})-B_{\ell}\right)+B_{\ell}$ (3.9)
$=W_{\ell}\,\tau_{W_{\ell},B_{\ell}}(X^{(\ell-1)})+B_{\ell},$
which proves (3.3). Going one step further,
$X^{(\ell)}$ $=\sigma(W_{\ell}X^{(\ell-1)}+B_{\ell})$
=$\stackrel{{\scriptstyle(*)}}{{=}}W_{\ell}W_{\ell-1}(W_{\ell}W_{\ell-1}){}^{+}\sigma(W_{\ell}(W_{\ell-1}\,\tau_{W_{\ell-1},b_{\ell-1}}(X^{(\ell-2)})+B_{\ell-1})+B_{\ell})$ (3.10)
where ()(*)( ∗ ) follows from the surjectivity of $W_{\ell}W_{\ell-1}$. An iteration of these arguments then yields (3.6). Finally,
$X^{(L)}$ $=W_{L}X^{(L-1)}+B_{L}$
$=W_{L}(W^{(L-1)}X^{(\tau,L-1)}+B^{(L-1)})+B_{L}$ (3.11)
$=W^{(L)}X^{(\tau,L-1)}+B^{(L)}.$
This lemma, and especially expression (3.7), suggest that we think of a neural network with $L$ layers as a concatenation of $L-1$ truncation maps acting on the data in input space $\mathbb{R}^{d_{0}}$, followed by an affine map to the appropriate output space $\mathbb{R}^{d_{L}}$. These truncation maps and the affine map are given by cumulative parameters, so that from now on in this paper a layer will be defined by prescribing $(W^{(\ell)}$, $b^{(\ell)})$ instead of $(W_{\ell},b_{\ell})$, for all $\ell=1\cdots,L$.
Lemma 3.3.
Consider a collection $\left\{W_{\ell}\right\}_{\ell=1}^{L}$ of surjective matrices, and assume the cumulative matrices $\{W^{(\ell)}\}_{\ell=1}^{L}$ defined as in (3.4) are also surjective. Define
$=(W^{(\ell)}){}^{+}W^{(\ell)}P^{(\ell-1)}.$ (3.12)
Then
$P^{(\ell)}=(W^{(\ell)}){}^{+}W^{(\ell)},$ (3.13)
and
$W^{(\ell)}P^{(\ell-1)}=W^{(\ell)}.$ (3.14)
Proof.
First, note that $P^{(1)}=(W^{(1)}){}^{+}W^{(1)}$ by definition, and assume $P^{(\ell-1)}=(W^{(\ell-1)}){}^{+}W^{(\ell-1)}$. If $W^{(\ell-1)}$ is surjective, then $W^{(\ell-1)}(W^{(\ell-1)}){}^{+}={\textbf{1}}$ and
$P^{(\ell)}$ $=(W^{(\ell)}){}^{+}W^{(\ell)}P^{(\ell-1)}$
$=(W^{(\ell)}){}^{+}W^{(\ell)}(W^{(\ell-1)}){}^{+}W^{(\ell-1)}$
$=(W^{(\ell)}){}^{+}W_{\ell}\left(W^{(\ell-1)}(W^{(\ell-1)}){}^{+}\right)W^{(\ell-1)}$ (3.15)
$=(W^{(\ell)}){}^{+}W_{\ell}W^{(\ell-1)}=(W^{(\ell)}){}^{+}W^{(\ell)},$
which gives (3.13).
$W^{(\ell)}({\textbf{1}}-P^{(\ell-1)})$ $=W^{(\ell)}Q^{(\ell-1)}$
$=W_{\ell}W^{(\ell-1)}Q^{(\ell-1)}$ (3.16)
$=0,$
and we have proved (3.14). ∎
$\mathbb{R}_{+}^{n}:=\left\{(x_{1},\cdots,x_{n})^{T}\in\mathbb{R}^{n}:x_{i}\geq 0,i=1,\cdots,n\right\}$ (3.17)
$y_{1}$
$y_{2}$
$x_{1}$
$x_{2}$
$Wx_{2}+b$
$Wx_{1}+b$
$W^{-1}(y_{2}-b)$
$W^{-1}(y_{1}-b)$
Figure 2. The action of $\tau_{W,b}$ in two dimensions. The image on the left represents the input space given by coordinates $(x_{1},x_{2})$. On the right we see the image of this space under the map $x\mapsto\sigma(Wx+b)$, where $\sigma$ acts according to the new coordinate system $(y_{1},y_{2})$: it is the identity on the positive sector (shaded in green), and can be thought of as projecting along the orange arrows in the other regions, which gives the total effect of mapping the negative sector (shaded in orange) to the origin. Pulling this action back to input space produces the cone on the picture to the left.
$\mathfrak{C}_{\theta}[h]:=\left\{x\in\mathbb{R}^{n}:\angle(x,h)\leq\frac{\theta}{2}\right\}.$ (3.18)
$\theta_{n}:=2\arccos\frac{\sqrt{n-1}}{\sqrt{n}},$ (3.19)
we note that $\mathfrak{C}_{\theta_{n}}[u_{n}]\subset\mathbb{R}^{n}$ is the largest cone centered around the diagonal $u_{n}=(1,\cdots,1)^{T}$ that is contained fully in $\mathbb{R}_{+}^{n}$. [^5] See appendix.. Similarly, the backward cone is fully contained in $\mathbb{R}_{-}^{n}$. A cone $\mathfrak{C}_{\theta}[u_{n}]$ of larger angle can be made to fit into $\mathbb{R}_{+}^{n}$ by a transformation which fixes the axis $u_{n}$ and shrinks every other direction perpendicular to the axis according to
$\lambda(\theta,\theta_{n}):=\begin{cases}\frac{\tan\frac{\theta_{n}}{2}}{\tan\frac{\theta}{2}}<1,&\text{ if }\theta>\theta_{n},\\ 1,&\text{ otherwise.}\end{cases}$ (3.20)
Now, in the case where there is no change in dimension, Lemma 3.4 below states that for any given cone we can construct $W$ and $b$ so that the truncation map $\tau_{W,b}$ acts as the identity on the forward cone, and projects the entirety of the backward cone to its base point. However, if $\tau_{W,b}$ is encoding a layer in a network for which there is a reduction in dimension, its image must also undergo a dimensional reduction. To account for the possible decrease in width of the network, consider the family of projectors
$\mathcal{I}(n,m)$ $:\mathbb{R}^{n}\to\mathbb{R}^{m},\quad n\geq m,$
$\mathcal{I}(n,m)(e^{n}_{i})$ =$=\begin{cases}e^{m}_{i},\quad i=1,\cdots,m,\\ 0,\quad\text{else,}\end{cases}$ (3.21)
where $\left\{e_{i}^{n}\right\}_{i=1}^{n}$ is the standard basis for $\mathbb{R}^{n}$, so that the matrix $\mathcal{I}(n,m)$ is a block matrix consisting of an identity matrix ${\textbf{1}}_{m\times m}$ and zeroes.
Lemma 3.4.
Consider the cone $p+\mathfrak{C}_{\theta}[h]\subset\mathbb{R}^{n}$ centered around a unit vector $h$ and based at $p$, and $W:\mathbb{R}^{n}\to\mathbb{R}^{m},\,b\in\mathbb{R}^{m}$ such that
$W$ $=\mathcal{I}(n,m)W_{\theta}R,$
$b$ $=-Wp,$ (3.22)
where
$W_{\theta}=\tilde{R}\,\operatorname{diag}(1,\lambda(\theta,\theta_{n}),\cdots,\lambda(\theta,\theta_{n}))\,\tilde{R}^{T},$ (3.23)
with $\lambda(\theta,\theta_{n})$ defined as in (3.20) and $R,\tilde{R}\in O(n,\mathbb{R})$ are such that
$Rh=\frac{u_{n}}{\left\lvert{u_{n}}\right\rvert},\,\,\tilde{R}e^{n}_{1}=\frac{u_{n}}{\left\lvert{u_{n}}\right\rvert}.$ (3.24)
Then
$\tau_{W,b}(x)=\begin{cases}((W){}^{+}W)\,x,\quad x\in p+\mathfrak{C}_{\theta}[h],\\ ((W){}^{+}W)\,p,\quad x\in p+\mathfrak{C}_{\theta}[-h].\end{cases}$ (3.25)
Moreover, if $P:\mathbb{R}^{n}\to\mathbb{R}^{n}$ is a map such that $WP=W$, then
$\tau_{W,b}(Px)=\tau_{W,b}(x).$ (3.26)
Proof.
First, consider $x\in p+\mathfrak{C}_{\theta}[h]$. Since
$\tau_{W,b}(x)=(W){}^{+}\left(\sigma(Wx+b)-b\right),$ (3.27)
$Wx+b$ $=W(x-p)=\mathcal{I}(n,m)W_{\theta}R(x-p),$ (3.28)
$R(x-p)\in\mathfrak{C}_{\theta}[Rh]=\mathfrak{C}_{\theta}[u_{n}].$ (3.29)
$W_{\theta}R(\mathfrak{C}_{\theta}[h])\subset\mathfrak{C}_{\theta_{n}}[u_{n}]\subset\mathbb{R}_{+}^{n},$ (3.30)
and $W(x-p)\in\mathbb{R}_{+}^{m}$.
$W(\mathfrak{C}_{\theta}[-h])\subset\mathfrak{C}_{\theta_{n}}[-u_{n}]\subset\mathbb{R}_{-}^{n},$ (3.31)
and then $W(x-p)\in\mathbb{R}_{-}^{m}$ implies $\sigma(Wx+b)=0$ and
$\tau_{W,b}(x)=(W){}^{+}(-b)=(W){}^{+}Wp.$ (3.32)
Finally, (3.26) follows easily from the definition of the truncation map,
$\tau_{W,b}(Px)$ $=(W){}^{+}\left(\sigma(W(Px)+b)-b\right)$
$=(W){}^{+}\left(\sigma(Wx+b)-b\right)$ (3.33)
$=\tau_{W,b}(x).$
We can now use the truncation map and Lemma 3.4 to generalize the construction of zero loss global minima from [CE23a] in two different directions. First, in section 4, we keep the training data as sufficiently separated, small enough clusters, but allow for $M=d_{0}\geq d_{1}\geq\cdots\geq d_{L}=Q$. Then, in section 5, we hold the width of every layer but the last one constant, $d_{0}=d_{1}=\cdots=d_{L-1}$, but allow for $M>Q$ and for a more general distribution of the training data.
Remark 3.5.
Note that the decomposition
$W=W_{\theta}R$ (3.34)
coincides with the polar decomposition of a matrix
$W=\left\lvert{W}\right\rvert R$ (3.35)
$=\tilde{R}\operatorname{diag}(1,\cdots)^{2}\tilde{R}^{T}$ (3.36)
$=W_{\theta}^{2}.$
4. Clustered data and dimensional reduction
We assume in this section the same geometry on the data as in [CE23a], which we now briefly recall. Consider
$\delta:=\sup_{i,j}\left\lvert{\Delta x_{0,j,i}}\right\rvert,$ (4.1)
and assume that
$\delta<c_{0}\min_{j}\left\lvert{\overline{x_{0,j}}-\overline{x}}\right\rvert,$ (4.2)
for $c_{0}\in(0,\frac{1}{4})$. Consider also
$\overline{x}:=\frac{1}{Q}\sum_{j=1}^{Q}\overline{x_{0,j}},$ (4.3)
noting that this is not necessarily the average of all data points, and for $j=1,\cdots,Q$, define
$f_{j}:=\frac{\overline{x}-\overline{x_{0,j}}}{\left\lvert{\overline{x}-\overline{x_{0,j}}}\right\rvert}$ (4.4)
to be the unit vector pointing from $\overline{x_{0,j}}$ to $\overline{x}$. Lastly, for each $j=1,\cdots,Q$, let $\theta_{*,j}$ be the smallest angle such that
$\bigcup_{j^{\prime}\neq j}B_{4\delta}(\overline{x_{0,j^{\prime}}})\subset\overline{x_{0,j}}+\mathfrak{C}_{\theta_{*,j}}[f_{j}],$ (4.5)
and assume
$\max_{j}\theta_{*,j}<\pi.$ (4.6)
Proposition 4.1.
Consider a set of training data $\mathcal{X}_{0}=\bigcup_{j=1}^{Q}\mathcal{X}_{0,j}\subset\mathbb{R}^{M}$ separated into $Q$ classes corresponding to linearly independent labels $\{y_{j}\}_{j=1}^{Q}\subset\mathbb{R}^{Q}$. If conditions (4.2) and (4.6) are satisfied, then a neural network with ReLU activation function defined as in (2), with $L=Q+1$ layers, $d_{0}=M$, $d_{Q+1}=Q$, and $d_{\ell-1}\geq d_{\ell}\geq Q$ for all $\ell=1,\cdots,L$, attains
$\min_{(W_{i},b_{i})_{i=1}^{L}}\mathcal{C}[(W_{i},b_{i})_{i=1}^{L}]=0,$ (4.7)
with weights and biases given recursively by (3.4), (3.5) and the following cumulative parameters: for $\ell=1,\cdots,Q$,
$W^{(\ell)}$ $=\mathcal{I}(d_{0},d_{\ell})W_{\theta_{*}}R_{\ell},$
$b^{(\ell)}$ $=-W^{(\ell)}(\overline{x_{0,\ell}}+\mu_{\ell}f_{\ell}),$ (4.8)
for $\mathcal{I}(d_{0},d_{\ell})$ a projection, $W_{\theta_{*}}\in GL(d_{0},\mathbb{R})$ and $R_{\ell}\in O(d_{0},\mathbb{R})$ constructed as in section 3, and some $\mu_{\ell}\in(2\delta,3\delta)$; for $\ell=Q+1$,
$W^{(Q+1)}$ $=Y(\overline{X_{0}^{red}}^{(\tau,Q)}){}^{+},$
$b^{(Q+1)}$ $=0.$ (4.9)
This minimum is degenerate.
Corollary 4.2.
To prove the proposition, we need to use barycentric coordinates. We consider a point $x\in\mathbb{R}^{M}$ to be represented by
$x=\sum_{j=1}^{Q}\kappa_{j}\overline{x_{0,j}}+\tilde{x}$ (4.10)
where $\tilde{x}\in(\operatorname{span}\{\overline{x_{0,j}}\}_{j=1}^{Q})^{\perp}$, and the means $\overline{x_{0,j}}$ form the vertices of a $Q$-simplex for which $\overline{x}$ is the barycenter. Switching to these coordinates can be done when preprocessing the data.
We defer the full proof to section 6.
5. Hyperplanes and sequentially separable data
In the previous section, we assumed that the data had to be neatly contained in sufficiently small and well spaced balls. However, when using Lemma 3.4 in the proof, it was sufficient to check that each ball lay inside a certain backward or forward cone. More generally, for the action of any given layer described by a truncation map, all that matters is that there exists a cone such that the class that is being singled out sits entirely within the backward cone, and the remaining data sits in the forward cone.
Definition 5.1.
We say a dataset is sequentially linearly separable if there exists an ordering of the classes such that for each $j=1,\cdots,Q$, there exists a point $p_{j}$ in the line segment $\{\overline{x}+(1-t)\overline{x_{0,j}}$, $t\in(0,1)\}$ and a hyperplane passing through $p_{j}$ that separates $\mathcal{X}_{0,j}$ and $\bigcup_{i<j}\{p_{i}\}\cup\bigcup_{i>j}\mathcal{X}_{0,i}$.

### v3

@@ Abstract
Keywords: deep learning, classification, loss landscape, interpolation, interpretability, geometry
1 Introduction
Deep learning algorithms are often thought to involve a trade-off between accuracy and interpretability. Highly overparameterized neural networks are capable of interpolating generic data, and gradient descent offers a standard way of training the parameters to minimize some loss function. However, the resulting models function as black boxes, providing limited insight into their internal mechanisms. In contrast, underparameterized networks cannot achieve zero training loss unless the data possesses sufficient structure (Chen and Ewald, 2025; Chen and Moore, 2025).
In this work, continuing the program started by the authors (Chen and Ewald, 2023a, b), we provide an interpretation for the action of the layers of a neural network, and explicitly construct global minima for a classification task with well distributed data, in a manner that is independent of the number of training samples. This characterization is done in terms of a map which encodes the action of a layer given by a weight matrix $W$, a bias vector $b$, and activation function $\sigma$,
$\tau_{W,b}(x)=(W){}^{+}\left(\sigma(Wx+b)-b\right),$
where $(W){}^{+}$ denotes the generalized inverse of $W$. A neural network with $L$ layers can be seen as an application of $L-1$ such transformations $\tau$ on input space, which curate the data, followed by an affine function in the last layer mapping to the output space,
$x^{(L)}$ $=W^{(L)}(x)^{(\tau,L-1)}+b^{(L)},$
where $x^{(\tau,L-1)}$ represents the $L-1$ consecutive applications of $\tau$ maps to an input $x$, and $W^{(L)},b^{(L)}$ are appropriate cumulative parameters. While this works for any given activation function, we specialize to the case of a ReLU neural network, and in this context call $\tau$ the truncation map. An incomplete, but highly intuitive visualization of its action comes from looking at cones: For any cone in input space, there exist $W$ and $b$ such that $\tau_{W,b}$ acts as the identity on the forward cone, and projects the entirety of the backward cone to the base point (see Lemma 2).
@@ 1 Introduction
Figure 1: A schematic representation of the action of a layer given by a truncation map. On the left, the original data with three classes contained in balls around the means $\overline{x_{0,j}}$, with $\mathcal{X}_{0,1}$ sitting in the backward cone $\mathfrak{C}_{\theta}$, and the remaining classes sitting in the forward cone. On the right, the truncation map has pushed all of $\mathcal{X}_{0,1}$ to a single point, the base of the cone. In green, we see the 2-simplex generated by the means, and the barycenter $\overline{x}$. The geometry of the data corresponds to the assumptions of Proposition 3.
Consider data in $\mathbb{R}^{M}$ to be classified into $Q$ classes. In previous work (Chen and Ewald, 2023a), the tools outlined above were used to construct explicit global minimizers for a network with constant width $M=Q$ and “clustered” data. In this paper, we extend this to the case where the width of the hidden layers is allowed to decrease monotonically from $M$ to $Q$, for $Q\leq M$. As a corollary, we show a reduction in the total number of parameters needed to interpolate the data. Our first result (informally stated) is
Theorem (Corollary to Proposition 3).
@@
As the angular aperture of a cone gets larger, it becomes a hyperplane. These are a well known classification tool, and there are algorithms dedicated to finding hyperplanes maximizing the distance to the two sets being separated, such as support vector machines (SVMs). Recent works show that under some circumstances, ReLU networks trained with gradient descent to perform binary classification tend to find max-margin hyperplanes (see e.g., Phuong and Lampert, 2020). [^1] The tendency of gradient-based algorithms to find critical points which generalize well or are simple in some way, e.g. low-rank, is called implicit bias or implicit regularization.